# Sequential Large Language Model-Based Hyperparameter Optimization (SLLMBO)

### v2

@@ Abstract
This study introduces SLLMBO, an innovative framework that leverages Large Language Models (LLMs) for hyperparameter optimization (HPO), incorporating dynamic search space adaptability, enhanced parameter landscape exploitation, and a hybrid, novel LLM-Tree-structured Parzen Estimator (LLM-TPE) sampler. By addressing limitations in recent fully LLM-based methods and traditional Bayesian Optimization (BO), SLLMBO achieves more robust optimization. This comprehensive benchmarking evaluates multiple LLMs—including GPT-3.5-turbo, GPT-4o, Claude-Sonnet-3.5, and Gemini-1.5-flash, extending prior work beyond GPT-3.5 and GPT-4 and establishing SLLMBO as the first framework to benchmark a diverse set of LLMs for HPO. By integrating LLMs’ established strengths in parameter initialization with the exploitation abilities demonstrated in this study, alongside TPE’s exploration capabilities, the LLM-TPE sampler achieves a balanced exploration-exploitation trade-off, reduces API costs, and mitigates premature early stoppings for more effective parameter searches. Across 14 tabular tasks in classification and regression, the LLM-TPE sampler outperformed fully LLM-based methods and achieved superior results over BO methods in 9 tasks. Testing early stopping in budget-constrained scenarios demonstrated competitive performance, indicating that LLM-based methods generally benefit from extended iterations for optimal results. This work lays the foundation for future research exploring open-source LLMs, reproducibility of LLM results in HPO, and benchmarking SLLMBO on complex datasets, such as image classification, segmentation, and machine translation.
Keywords: Large Language Models, Hyperparameter Optimization, Bayesian Optimization
@@ 1.2 Limitations of Bayesian Optimization
To overcome the limitations of bayesian optimization, various ideas have been proposed, such as transfer-learned-based search spaces to reduce human effort (Shahriari et al., 2016; Nguyen et al., 2019); designing algorithmic adaptive search spaces to solve fixed search space problems (Perrone et al., 2019; Li et al., 2022); multi-fidelity optimization (Hutter et al., 2019; Joy et al., 2020) or early-stopping (Tan et al., 2024) to speed up sequential learning; warm-starting to initialize parameters based on successful parameter values learned on similar datasets (Feurer et al., 2014; Kim et al., 2017; Bai et al., 2023; Poloczek et al., 2016). However, each method suits a specific problem and does not lead to a unified framework.
@@ 1.5 Proposed Approach: Sequential Large Language Model-Based Optimization (SLLMBO)
To address the identified research gaps, we propose SLLMBO, a novel method that leverages the strengths of LLMs and combines them with tree-structured parzen estimator sampling. SLLMBO features an adaptive search space that evolves based on prior iterations and employs a hybrid LLM-TPE Sampler. It warm-starts with LLM and alternates between LLM-based sampling and TPE-based exploration to balance exploration and exploitation. Additionally, early stopping is integrated to terminate the optimization when no improvements are observed, thus reducing unnecessary computational costs.
@@ 2.2 Bayesian Optimization
To overcome the limitations of GPs, alternative surrogate models like the tree-structured parzen estimator (TPE) (Bergstra et al., 2011) and sequential model-based algorithm configuration (SMAC) (Hutter et al., 2011) have been developed. TPE models the objective function by dividing the search space into two distributions with a predefined threshold $y^{*}$, as expressed in Equation 2:
@@ 3.6 Reasoning-Enabled Optimization
To enhance transparency and interpretability, SLLMBO offers an optional Reasoning Engine. This feature requires the LLM to explain each decision made during the optimization process: whether to update or retain the current search space and why specific parameters were chosen. These explanations are stored alongside the optimization history, allowing for post-hoc analysis of the optimization decisions.
@@ 3.8 Experimental Setup
The experimental evaluation of SLLMBO was conducted on six datasets: three classification and two regression tasks from The UCI Machine Learning Repository (Kelly et al., ), one regression task from OpenML (Vanschoren et al., 2014), and the M5 Forecasting Accuracy dataset (Howard et al., 2020), a challenging multiple time series forecasting task. The classification datasets include gas_drift (Vergara, 2012), cover_type (Blackard, 1998), and adult_census (Becker and Kohavi, 1996), while the regression datasets include bike_sharing (Fanaee-T, 2013), concrete strength (Vanschoren et al., 2014), and energy (Candanedo, 2017). For each dataset, approximately 20% of the data is kept for testing the performance of the best parameters found by some HPO method on the rest of the data, helping to explore overfitting. To assess the performance of each iteration during the HPO process, the F1-score was used for classification tasks, while the mean absolute error (MAE) was employed for regression tasks. For parameter suggestion in subsequent iterations, the average cross-validation score across the folds was used as the performance metric. Detailed descriptions of each dataset, including feature and target column selection, preprocessing, feature engineering, and the specific cross-validation techniques applied, can be found in Appendix B.
To evaluate the performance of SLLMBO, it was compared against state-of-the-art Bayesian Optimization methods using two popular libraries: Optuna (Akiba et al., 2019) and Hyperopt (Bergstra et al., 2015), both leveraging the Tree-structured Parzen Estimator (TPE) as their sampling strategy. The choice of Optuna and Hyperopt is grounded in recent studies highlighting their effectiveness and widespread use in tabular tasks (Hassanali et al., 2024; Lai et al., 2023; Lv, 2023; Arden and Safitri, 2022; Putatunda and Rama, 2019; Wang and Wang, 2020; Omkari and Shinde, 2022). LightGBM and XGBoost were used as machine learning models, known for their superior performance on tabular datasets (Lai et al., 2023; Lv, 2023; Arden and Safitri, 2022; Wang and Wang, 2020; Omkari and Shinde, 2022).
@@ 4.1 Initial Exploration: SLLMBO with Fully LLM-powered Methods and Intelligent Summary
Figure 2: The Figure illustrates the optimization history for initial LLM HPO strategy with GPT-3.5-Turbo and intelligent_summary, Optuna, and Hyperopt strategies and LLM’s early stopping. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM
@@
Despite these challenges, early stopping proved advantageous in some instances, such as in the energy dataset with XGBoost and the M5 dataset with LightGBM, where the LLM-based method achieved superior or comparable results using far fewer iterations. In these cases, the LLM method required, on average, 6 times less computational budget than Optuna and 8.5 times less than Hyperopt while still idenpngying better or near-optimal parameters (see Table 2). However, while overexploitation worked effectively when the initial parameters were near the global optimum, it limited the search’s effectiveness when the initial parameters were in the vicinity of a local optimum. This highlights the need for a better balance between exploration and exploitation, particularly when the starting point is less ideal.
@@
As the analysis progressed, it became clear that early stopping was not solely due to the LLM’s optimization behavior. Still, it was also influenced by external factors such as API-related issues and LLM-based history summarization. These technical limitations highlighted the need for more robust memory management and request handling. To address these technical concerns, the decision was made to transition to the LangChain framework, which offers automatic memory management and improved error-handling mechanisms, mitigating the issues encountered in the earlier experiments. While this switch was not explicitly aimed at solving the overexploitation problem, LangChain’s more stable and controlled environment allowed for further analysis of the implicit effects on the exploration-exploitation trade-off. This transition also allowed for easy experimentation of other LLM models, such as GPT-4o, Gemini, and Claude-Sonnet, to assess whether these models could improve the optimization process. Furthermore, in the next phase of experiments, new sampling strategies specifically designed to address overexploitation will be discussed in detail using LangChain.
@@ 4.2 SLLMBO with Fully LLM-Powered Methods and LangChain
Figure 4: The Figure illustrates the optimization history for LLMs: GPT-4o, GPT-3.5-Turbo, Claude-3.5-Sonnet, and Gemini-1.5-Flash with second LLM strategy, using LangChain to run the LLM APIs, replacing intelligent summary with LangChain’s memory buffer. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@ 4.3 SLLMBO with LLM-TPE Sampler
To achieve this, GPT-4o and Gemini-1.5-Flash were selected. Gemini was the fastest and cheapest model, but it had premature stoppings, so it is included to assess whether the hybrid strategy could improve its performance. GPT-3.5-turbo was excluded due to its consistent success in previous phases, and Sonnet was omitted due to high costs.
@@
To further assess the effectiveness of LLM-based initialization in hyperparameter optimization, the GPT-4o-based LLM-TPE sampler was tested with random and LLM initialization strategies. The results in Tables 5 and 6 reaffirmed the strength of LLM-based initialization, as in 13 out of the 14 tasks, GPT-4o’s LLM initialization outperformed random initialization in providing a better starting point. However, a more detailed analysis of the final optimized parameters after 50 iterations revealed a notable distinction between regression and classification tasks. In regression tasks, as given in Table 6, the best parameters starting with random initialization consistently resulted in better test scores. In contrast, Table 5 shows that, in nearly all classification tasks, starting with GPT-4o initialization led to superior results.
@@
Figure 5: Optimization history plots for LLM-TPE Sampler with LLMs GPT-4o and Gemini-1.5-Flash with LLM-based initialization, also GPT-4o with random initialization. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@
The success of the LLM-TPE sampler can be attributed to its ability to maintain a balanced exploration-exploitation trade-off. The fully LLM-based approach excelled at exploitation, as results from Subsections 4.1 and 4.2 demonstrated. At the same time, the TPE method’s capacity to explore uncertain regions of parameter space is well-established in prior research. By combining these two approaches, the LLM-TPE sampler offered a more holistic and effective optimization strategy, addressing the limitations of using either technique in isolation. This integration significantly reduced the occurrence of premature early stoppings observed in previous methods. As shown from the optimization history plots in Figure 5, the LLM-TPE sampler enabled models to complete 50 iterations with fewer API calls, reducing computational costs. Furthermore, the analysis of optimization scores throughout the trials revealed that the LLM-TPE sampler no longer exhibited the limited exploration and narrow score ranges characteristic of fully LLM-based strategies. Instead, it achieved a more balanced exploration, with score variability comparable to Bayesian Optimization methods. This enhanced exploration facilitated the discovery of more optimal parameter configurations, contributing to the improved performance observed across multiple tasks.
To evaluate the effectiveness of the LLM-TPE sampler in budget-constrained scenarios, early stopping (ES) with a patience of 5 is tested, too. This experiment focused on the Gemini-1.5-Flash LLM-TPE sampler, comparing its performance with and without early stopping and against Optuna to determine how limiting the number of iterations influences hyperparameter optimization outcomes. While early stopping reduced computational costs across tasks, it also affected performance (Tables 78). Gemini with ES outperformed Optuna with ES in 6 tasks, yet both showed decreased performance compared to their non-ES counterparts. Optuna benefitted from early stopping in specific tasks, while Gemini exhibited a consistent performance drop.
Figure 6: Optimization history plots for LLM-TPE Sampler with Gemini-1.5-Flash with LLM initialization and Optuna with patience 5. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@ 5 Conclusion
This research explored the potential of large language models (LLMs) for hyperparameter optimization by comparing various models: GPT-3.5-Turbo, GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Flash, with traditional bayesian optimization methods such as optuna and hyperopt. The study comprehensively analyzed LLMs’ performance in parameter initialization and optimization, adaptive search spaces, and the management of exploration-exploitation trade-offs. Two key contributions emerged: a novel hybrid LLM-TPE sampler for improved exploration-exploitation balance and an adaptive search space mechanism driven by LLMs to automatically evolve the parameter space. Early stopping criteria were also integrated to reduce computational costs. Initial experiments with fully LLM-powered methods demonstrated that LLMs consistently outperformed random initialization, particularly in regression tasks, and offered competitive performance in hyperparameter optimization compared to traditional BO methods. However, challenges such as overexploitation and limited exploration arose, particularly with complex datasets. To address these issues, the introduction of LangChain solved technical challenges, including memory management and request handling. It enabled longer iterations and improved performance in parameter initialization and hyperparameter tuning. While LangChain enhanced stability and reduced early stoppings, it did not entirely eliminate the problem of overexploitation. The LLM-TPE hybrid sampler combined the LLMs’ strength in initialization and exploitation with TPE’s exploration capabilities, resulting in a significantly better balance between exploration and exploitation. This method delivered superior HPO performance across multiple tasks compared to fully LLM-based and traditional BO methods. It also reduced premature early stopping and minimized computational costs through fewer LLM API calls. Testing the LLM-TPE sampler under budget-constrained scenarios with early stopping revealed that while it remained competitive, LLM-based methods generally required more iterations to reach optimal results when subjected to strict early stopping criteria. While this study provided a comprehensive analysis of various API-based LLMs in hyperparameter optimization, future work could extend this research by exploring the performance of open-source LLMs. This would offer broader accessibility and potentially reduce the reliance on paid APIs. Furthermore, this study focused on tabular datasets due to API-related cost constraints. Future investigations could apply the proposed methods to more complex tasks, such as image classification, segmentation, object detection, or machine translation, to evaluate their effectiveness across a broader range of domains. Lastly, it is essential to note that despite the promising results, LLMs face a significant limitation: their lack of reproducibility, as there is no guarantee of retrieving identical optimization histories across runs. The complete codebase for this study is available at https://github.com/KananMahammadli/SLLMBO.

### v3

# Sequential Large Language Model-Based Hyper-parameter Optimization

@@ Abstract
This study introduces SLLMBO, an innovative framework leveraging large language models (LLMs) for hyperparameter optimization (HPO), incorporating dynamic search space adaptability, enhanced parameter space exploitation, and a novel LLM-tree-structured parzen estimator (LLM-TPE) sampler. By addressing limitations in recent fully LLM-based methods and traditional bayesian optimization (BO), SLLMBO achieves more robust optimization. This comprehensive benchmarking evaluates multiple LLMs—including GPT-3.5-Turbo, GPT-4o, Claude-Sonnet-3.5, and Gemini-1.5-Flash, extending prior work and establishing SLLMBO as the first framework to benchmark a diverse set of LLMs for HPO. By integrating LLMs’ established strengths in parameter initialization with the exploitation abilities demonstrated in this study, alongside TPE’s exploration capabilities, the LLM-TPE sampler achieves a balanced exploration-exploitation trade-off, reduces API costs, and mitigates premature early stoppings for more effective parameter searches. Across 14 tabular tasks in classification and regression, the LLM-TPE sampler outperformed fully LLM-based methods and achieved superior results over BO methods in 9 tasks. Testing early stopping in budget-constrained scenarios demonstrated competitive performance, indicating that LLM-based methods generally benefit from extended iterations for optimal results. This work lays the foundation for future research exploring open-source LLMs, reproducibility of LLM results in HPO, and benchmarking SLLMBO on complex datasets, such as image classification, segmentation, and machine translation.
Keywords: Large Language Models, Hyperparameter Optimization, Bayesian Optimization, Tree-Structured Parzen Estimator, Machine Learning
@@ 1.2 Limitations of Bayesian Optimization
Various ideas have been proposed to overcome the limitations of bayesian optimization, such as transfer-learned-based search spaces to reduce human effort (Shahriari et al., 2016; Nguyen et al., 2019); designing algorithmic adaptive search spaces to solve fixed search space problems (Perrone et al., 2019; Li et al., 2022); multi-fidelity optimization (Hutter et al., 2019; Joy et al., 2020) or early-stopping (Tan et al., 2024) to speed up sequential learning; warm-starting to initialize parameters based on successful parameter values learned on similar datasets (Feurer et al., 2014; Kim et al., 2017; Bai et al., 2023; Poloczek et al., 2016). However, each method suits a specific problem and does not lead to a unified framework.
@@ 1.5 Proposed Approach: Sequential Large Language Model-Based Optimization (SLLMBO)
We propose SLLMBO, a novel method that leverages the strengths of LLMs and combines them with tree-structured parzen estimator sampling to address the identified research gaps. SLLMBO features an adaptive search space that evolves based on prior iterations and employs a hybrid LLM-TPE Sampler. It warm-starts with LLM and alternates between LLM-based sampling and TPE-based exploration to balance exploration and exploitation. Additionally, early stopping is integrated to terminate the optimization when no improvements are observed, thus reducing unnecessary computational costs.
@@ 2.2 Bayesian Optimization
Alternative surrogate models like the tree-structured parzen estimator (TPE) (Bergstra et al., 2011) and sequential model-based algorithm configuration (SMAC) (Hutter et al., 2011) have been developed to overcome the limitations of GPs. TPE models the objective function by dividing the search space into two distributions with a predefined threshold $y^{*}$, as expressed in Equation 2:
@@ 3.6 Reasoning-Enabled Optimization
SLLMBO offers an optional Reasoning Engine to enhance transparency and interpretability. This feature requires the LLM to explain each decision made during the optimization process: whether to update or retain the current search space and why specific parameters were chosen. These explanations are stored alongside the optimization history, allowing for post-hoc analysis of the optimization decisions.
@@ 3.8 Experimental Setup
The experimental evaluation of SLLMBO was conducted on six datasets: three classification and two regression tasks from The UCI Machine Learning Repository (Kelly et al., ), one regression task from OpenML (Vanschoren et al., 2014), and the M5 Forecasting Accuracy dataset (Howard et al., 2020), a challenging multiple time series forecasting task. The classification datasets include gas_drift (Vergara, 2012), cover_type (Blackard, 1998), and adult_census (Becker and Kohavi, 1996), while the regression datasets include bike_sharing (Fanaee-T, 2013), concrete strength (Vanschoren et al., 2014), and energy (Candanedo, 2017). For each dataset, approximately 20% of the data is kept for testing the performance of the best parameters found by some HPO method on the rest of the data, helping to explore overfitting. The F1-score was used for classification tasks to assess the performance of each iteration during the HPO process, while the mean absolute error (MAE) was employed for regression tasks. For parameter suggestion in subsequent iterations, the average cross-validation score across the folds was used as the performance metric. Detailed descriptions of each dataset, including feature and target column selection, preprocessing, feature engineering, and the specific cross-validation techniques applied, can be found in Appendix B.
The SLLMBO was compared against state-of-the-art Bayesian Optimization methods using two popular libraries: Optuna (Akiba et al., 2019) and Hyperopt (Bergstra et al., 2015), both leveraging the Tree-structured Parzen Estimator (TPE) as their sampling strategy. The choice of Optuna and Hyperopt is grounded in recent studies highlighting their effectiveness and widespread use in tabular tasks (Hassanali et al., 2024; Lai et al., 2023; Lv, 2023; Arden and Safitri, 2022; Putatunda and Rama, 2019; Wang and Wang, 2020; Omkari and Shinde, 2022). LightGBM and XGBoost were used as machine learning models, known for their superior performance on tabular datasets (Lai et al., 2023; Lv, 2023; Arden and Safitri, 2022; Wang and Wang, 2020; Omkari and Shinde, 2022).
@@ 4.1 Initial Exploration: SLLMBO with Fully LLM-powered Methods and Intelligent Summary
Figure 2: The Figure illustrates the optimization history for Optuna, Hyperopt, and initial LLM HPO strategy with GPT-3.5-Turbo and intelligent_summary and LLM’s early stopping with patience_15. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@
Despite these challenges, early stopping proved advantageous in some instances, such as in the energy dataset with XGBoost and the M5 dataset with LightGBM, where the LLM-based method achieved superior or comparable results using far fewer iterations. In these cases, the LLM method required, on average, 6 times less computational budget than Optuna and 8.5 times less than Hyperopt while still identifying better or near-optimal parameters (see Table 2). However, while overexploitation worked effectively when the initial parameters were near the global optimum, it limited the search’s effectiveness when the initial parameters were in the vicinity of a local optimum. This highlights the need for a better balance between exploration and exploitation, particularly when the starting point is less ideal.
@@
As the analysis progressed, it became clear that early stopping was not solely due to the LLM’s optimization behavior. Still, it was also influenced by external factors such as API-related issues and LLM-based history summarization. These technical limitations highlighted the need for more robust memory management and request handling. The decision was made to transition to the LangChain framework to address these technical concerns, which offers automatic memory management and improved error-handling mechanisms, mitigating the issues encountered in the earlier experiments. While this switch was not explicitly aimed at solving the overexploitation problem, LangChain’s more stable and controlled environment allowed for further analysis of the implicit effects on the exploration-exploitation trade-off. This transition also allowed for easy experimentation of other LLM models, such as GPT-4o, Gemini, and Claude-Sonnet, to assess whether these models could improve the optimization process. Furthermore, in the next phase of experiments, new sampling strategies specifically designed to address overexploitation will be discussed in detail using LangChain.
@@ 4.2 SLLMBO with Fully LLM-Powered Methods and LangChain
Figure 4: The Figure illustrates the optimization history for LLMs: GPT-4o, GPT-3.5-Turbo, Claude-3.5-Sonnet, and Gemini-1.5-Flash with second LLM strategy, using LangChain to run the LLM APIs, replacing intelligent summary with LangChain’s memory buffer. Early Stopping of patience_15 is used for all LLMs. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@ 4.3 SLLMBO with LLM-TPE Sampler
GPT-4o and Gemini-1.5-Flash were selected to achieve this goal. Gemini was the fastest and cheapest model, but it had premature stoppings, so it is included to assess whether the hybrid strategy could improve its performance. GPT-3.5-turbo was excluded due to its consistent success in previous phases, and Sonnet was omitted due to high costs.
@@
The GPT-4o-based LLM-TPE sampler was tested with random and LLM initialization strategies to further assess the effectiveness of LLM-based initialization in hyperparameter optimization. The results in Tables 5 and 6 reaffirmed the strength of LLM-based initialization, as in 13 out of the 14 tasks, GPT-4o’s LLM initialization outperformed random initialization in providing a better starting point. However, a more detailed analysis of the final optimized parameters after 50 iterations revealed a notable distinction between regression and classification tasks. In regression tasks, as given in Table 6, the best parameters starting with random initialization consistently resulted in better test scores. In contrast, Table 5 shows that, in nearly all classification tasks, starting with GPT-4o initialization led to superior results.
@@
Figure 5: Optimization history plots for LLM-TPE Sampler with GPT-4o with LLM-based initialization, random initialization, and Gemini-1.5-Flash with LLM initialization. Early Stopping of patience_15 is used for all LLMs. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@
The success of the LLM-TPE sampler can be attributed to its ability to maintain a balanced exploration-exploitation trade-off. The fully LLM-based approach excelled at exploitation, as results from Sections 4.1 and 4.2 demonstrated. At the same time, the TPE method’s capacity to explore uncertain regions of parameter space is well-established in prior research. By combining these two approaches, the LLM-TPE sampler offered a more holistic and effective optimization strategy, addressing the limitations of using either technique in isolation. This integration significantly reduced the occurrence of premature early stoppings observed in previous methods. As shown from the optimization history plots in Figure 5, the LLM-TPE sampler enabled models to complete 50 iterations with fewer API calls, reducing computational costs. Furthermore, the analysis of optimization scores throughout the trials revealed that the LLM-TPE sampler no longer exhibited the limited exploration and narrow score ranges characteristic of fully LLM-based strategies. Instead, it achieved a more balanced exploration, with score variability comparable to Bayesian Optimization methods. This enhanced exploration facilitated the discovery of more optimal parameter configurations, contributing to the improved performance observed across multiple tasks.
Early stopping (ES) with patience of 5 is also tested to evaluate the effectiveness of the LLM-TPE sampler in budget-constrained scenarios. This experiment focused on the Gemini-1.5-Flash LLM-TPE sampler, comparing its performance with and without early stopping and against Optuna to determine how limiting the number of iterations influences hyperparameter optimization outcomes. While early stopping reduced computational costs across tasks, it also affected performance (Tables 78). Gemini with ES outperformed Optuna with ES in 6 tasks, yet both showed decreased performance compared to their non-ES counterparts. Optuna benefitted from early stopping in specific tasks, while Gemini exhibited a consistent performance drop.
Figure 6: Optimization history plots for Optuna and LLM-TPE Sampler with Gemini-1.5-Flash with LLM initialization, both having Early Stopping of patience_5. The top panel represents the energy dataset with the XGBoost model, the middle panel bike sharing dataset with LightGBM, and the bottom panel cement strength dataset with LightGBM.
@@ 5 Conclusion
This research explored the potential of large language models (LLMs) for hyperparameter optimization by comparing various models: GPT-3.5-Turbo, GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Flash, with traditional bayesian optimization methods such as optuna and hyperopt. The study comprehensively analyzed LLMs’ performance in parameter initialization and optimization, adaptive search spaces, and the management of exploration-exploitation trade-offs. Two key contributions emerged: a novel hybrid LLM-TPE sampler for improved exploration-exploitation balance and an adaptive search space mechanism driven by LLMs to automatically evolve the parameter space. Early stopping criteria were also integrated to reduce computational costs. Initial experiments with fully LLM-powered methods demonstrated that LLMs consistently outperformed random initialization, particularly in regression tasks, and offered competitive performance in hyperparameter optimization compared to traditional BO methods. However, challenges such as overexploitation and limited exploration arose, particularly with complex datasets. The introduction of LangChain solved technical challenges, including memory management and request handling. It enabled longer iterations and improved performance in parameter initialization and hyperparameter tuning. While LangChain enhanced stability and reduced early stoppings, it did not entirely eliminate the problem of overexploitation. The LLM-TPE hybrid sampler combined the LLMs’ strength in initialization and exploitation with TPE’s exploration capabilities, resulting in a significantly better balance between exploration and exploitation. This method delivered superior HPO performance across multiple tasks compared to fully LLM-based and traditional BO methods. It also reduced premature early stopping and minimized computational costs through fewer LLM API calls. Testing the LLM-TPE sampler under budget-constrained scenarios with early stopping revealed that while it remained competitive, LLM-based methods generally required more iterations to reach optimal results when subjected to strict early stopping criteria. While this study provided a comprehensive analysis of various API-based LLMs in hyperparameter optimization, future work could extend this research by exploring the performance of open-source LLMs. This would offer broader accessibility and potentially reduce the reliance on paid APIs. Furthermore, this study focused on tabular datasets due to API-related cost constraints. Future investigations could apply the proposed methods to more complex tasks, such as image classification, segmentation, object detection, or machine translation, to evaluate their effectiveness across a broader range of domains. Lastly, it is essential to note that despite the promising results, LLMs face a significant limitation: their lack of reproducibility, as there is no guarantee of retrieving identical optimization histories across runs. The complete codebase for this study is available at https://github.com/KananMahammadli/SLLMBO.